# Probabilistic Robustness for Free? Revisiting Training via a Benchmark

### v1

@@ E.3 Analysis on RT Learning
Here, we provide a proof of the Lipschitz and smoothness conditions for the RT method. As shown in Wang et al. (2021), the optimization process follows gradient descent with perturbations sampled from the distribution $\Pr(\cdot\mid\bm{x})$, which can be interpreted as a form of data augmentation within standard training. Consequently, it shares the same theoretical generalization error bounds as standard training.
@@ E.4 Analysis on AT-PR
Alg. 2 essentially searches for the region $\Delta_{i}$ with the largest coverage along the loss surface. Two cases arise from this selection, which are also discussed in Zhang et al. (2025):
@@
In the first case, the generalization gap decreases (as shown in Thm. 1), which corresponds to the experimental results of WRN-28-10 on CIFAR-100. In the second case, the generalization gap of AT-PR is comparable to that of PGD, as shown in the CIFAR-10 experiments (See Table. 9). Our experiments, along with results from Zhang et al. (2025), demonstrate that approximately $90\%$ of inputs correspond to the second case. This corresponds to the special case where only a single AE candidate is considered in Alg. 2.
@@ E.5 Uniform Algorithm Stability Analysis on Adversarial Training
Theorem 5 (Generalization in Expectation by Hardt et al. (2016))

### v2

@@ E.3 Analysis on RT Learning
Here, we provide a proof of the Lipschitz and smoothness conditions for the RT method. As shown in wang2021statistically, the optimization process follows gradient descent with perturbations sampled from the distribution $\Pr(\cdot\mid\bm{x})$, which can be interpreted as a form of data augmentation within standard training. Consequently, it shares the same theoretical generalization error bounds as standard training.
@@ E.4 Analysis on AT-PR
Alg. 2 essentially searches for the region $\Delta_{i}$ with the largest coverage along the loss surface. Two cases arise from this selection, which are also discussed in zhang2025adversarial:
@@
In the first case, the generalization gap decreases (as shown in Thm. 1), which corresponds to the experimental results of WRN-28-10 on CIFAR-100. In the second case, the generalization gap of AT-PR is comparable to that of PGD, as shown in the CIFAR-10 experiments (See Table. 9). Our experiments, along with results from zhang2025adversarial, demonstrate that approximately $90\%$ of inputs correspond to the second case. This corresponds to the special case where only a single AE candidate is considered in Alg. 2.
@@ E.5 Uniform Algorithm Stability Analysis on Adversarial Training
Theorem 5 (Generalization in Expectation by hardt2016train)